### v2

# Robust $\mathcal{H}_{\infty}$ control under stochastic requirements: minimizing conditional value-at-risk instead of worst-case performance

Abstract
Conventional robust $\mathcal{H}_{2}/\mathcal{H}_{\infty}$ control minimizes the worst-case performance, often leading to a conservative design driven by very rare parametric configurations. To reduce this conservatism while taking advantage of the stochastic properties of Monte-Carlo sampling and its compatibility with parallel computing, we introduce an alternative paradigm that optimizes the controller with respect to a stochastic criterion, namely the conditional value at risk. We illustrate the potential of this approach on a realistic satellite benchmark, showing that it can significantly improve overall performance by tolerating some degradation in very rare worst-case scenarios.
I Introduction
Robust structured $\mathcal{H}_{\infty}$ control [5] traditionally aims at finding a controller that minimizes the worst-case performance across all possible parametric configurations:
[…]
where $T_{zw}(\mathrm{s},k,\delta)$ is a transfer function between input $w$ and output $z$, $k$ is a vector of decision variables defined over an admissible space $\mathcal{K}$ and describing a controller $K$ of chosen structure (fixed-order, PID, etc.), and $\delta$ is a vector of uncertain parameters. In this formulation, the space $\mathcal{D}_{\delta}$ of uncertain parameters is deterministic, and is generally the hypercube $1,1]^{m}$ after normalization of the $m$ uncertain parameters. The case of nonlinear constraint restricting this parametric space was also addressed in [15]. When there are multiple control objectives, it is sought to solve:
[…]
where $||\cdot||$ may refer to the $\mathcal{H}_{2}$ or $\mathcal{H}_{\infty}$ system norm. The channel $w_{1}\rightarrow z_{1}$ corresponds to a soft requirement to be minimized, while the other channels $w_{i}\rightarrow z_{i}$ ($i>1$) are hard requirements, that must be inferior to 1 (once normalized) but not necessarily minimized. As this formulation does not involve any notion of probability, the worst-case configuration is often extremely unlikely, leading to a conservative design with regard to industrial requirements that are generally defined in a stochastic sense and relying on a probabilistic description of the uncertain parameters (e.g. gaussian, uniform, or other distributions).
When considering deterministic parameters $\delta\in[-1,1]^{m}$, the state-of-art for robust $\mathcal{H}_{\infty}$ control is the nonsmooth optimization algorithm of [2], implemented in MATLAB’s function systune, that solves the inner maximization problem of (1) using a local search enabled by the computation of Clarke subdifferentials [11] of the $\mathcal{H}_{\infty}$ norm [9]. Identified worst-case configurations are iteratively added to an active set $\mathcal{D}_{a}\subset\mathcal{D}$ on which the outer minimization problem is solved based on the algorithm of [4]. Then, global robustness verification can be done for example with Monte-Carlo sampling [27] or with $\mu$-analysis [12]. The former provides statistical guarantees based on the worst-case sample, while the latter, generally based on branch-and-bound algorithms [21], provides a formal guarantee of detecting the global worst-case configuration. However, as noted above, this worst-case scenario can be extremely rare, which motivated the development of probabilistic $\mu$-analysis [8]. Nevertheless, as most global optimization methods, $\mu$-based techniques exhibit limited scalability with the number of parameters; in fact, they attempt to solve a NP-hard problem [28]. Furthermore, they are not oriented toward robust controller synthesis, which is the primary concern of the present work.
This paper proposes an alternative paradigm to perform robust $\mathcal{H}_{2}/\mathcal{H}_{\infty}$ controller synthesis, with two objectives: (i) to optimize a robust controller in the sense of stochastic requirements, thereby ensuring a better match between the control synthesis framework and the stochastic model which better reflects the true underlying uncertainty-generating process than the deterministic formulation, thus reducing conservatism, and (ii) to leverage Monte-Carlo methods, which are commonly used in stochastic programming [23], come with stochastic guarantees, and naturally exploit parallel computing.
For this, we rely on the conditional value at risk (CVAR), a common concept in the field of finances to quantify potential extreme losses in a portfolio. The main idea of this approach is to consider $T_{zw}(\mathrm{s},k,\delta)||$ as a random variable defined over a probabilistic space rather than a deterministic function defined over a hypercube. This formulation is introduced in Section II, and adapted in Section III to be more suitable for optimization. Section IV-A proposes a solution method and identifies open challenges. Finally, Section V illustrates this method on the robust control of a mechanical system.
We note that the use of CVAR in control problems is not new; see, for example, [13, 16, 29]. However, to the best of the authors’ knowledge, its application to the robust $\mathcal{H}_{\infty}$ control problem, traditionally treated via the worst-case approach described above, has not been previously discussed.
II Problem statement
Considering a vector $\delta\in\ \mathbb{R}^{m}$ of uncertain parameters and a controller $K$ structured by a vector $k\in\mathcal{K}$ of decision variables (e.g. fixed-order controller, observer-based, PID, etc.), let us consider a transfer function $T_{zw}(\mathrm{s},k,\delta)$ under the Linear Fractional Representation (LFR) as in Fig. 1, where $\Delta=\textrm{diag}(\delta_{i}I_{n_{i}})$ ($n_{i}$ is the number of repetitions of $\delta_{i}$).
Figure 1: LFR representation of $T_{zw}(\mathrm{s},k,\delta)$
Assumption 1
The random variable $\delta$ has bounded support $\mathcal{D}_{\delta}$, and its probability distribution admits a density $p(\delta)$.
Assuming a bounded support is common in control engineering, as it is necessary in the traditional deterministic framework to define a worst case (as in Eq. (1)). In this paper, it is necessary to realistically make Assumption 2 below. Without loss of generality, we consider that $\mathcal{D}_{\delta}\subset[-1,1]^{m}$ and that $\delta=0$ corresponds to the nominal configuration.
For a system norm $||\cdot||$ which may be the $\mathcal{H}_{2}$ or $\mathcal{H}_{\infty}$ norm [^1] For the $\mathcal{H}_{2}$ norm, we assume that the matrix $D$ of the state-space representation is such the $\mathcal{H}_{2}$ norm is correctly defined, cf. e.g. [19, 3, 6]. in this paper, the loss function:
[…]
is a random variable that depends on the decision vector $k$.
Assumption 2
Given $k\in\mathcal{K}$, $L_{zw}(k,\delta)$ satisfies Assumption 2 if $T_{zw}(\mathrm{s},k,\delta)$ is internally stable for all $\delta\in\mathcal{D}_{\delta}$.
Assumption 2 allows the loss function to be well defined for any $\delta$, and thus to define the expected value in the following. This assumption might be seen as restrictive; but in practice, it can anyway be desirable to guarantee stability even in rare worst-case scenarios, with the following stochastic formulation relaxing only the performance requirements measured with the $\mathcal{H}_{\infty}$ or $\mathcal{H}_{2}$ norms.
Assumption 3
Given $k\in\mathcal{K}$, $L_{zw}(k,\delta)$ satisfies Assumption 3 when its distribution function is continuous.
Assumption 3 is adopted here following the work of [20] to properly introduce Definition 1.
Definition 1 ($\beta$-VAR, $\beta$-CVAR)
Given the loss function $L_{zw}(k,\delta)$ satisfying Assumptions 1, 2 and 3, the $\beta$-Value at risk ($\beta$-VAR), noted $\text{VAR}_{\beta}^{zw}(k)$, is:
$\text{VAR}_{\beta}^{zw}(k)=\text{min}\{\alpha\in\mathbb{R}:\Psi^{zw}(k,\alpha)\geq\beta\}\;.$ (4)
where $\Psi^{zw}(k,\alpha)$ is the probability of $L_{zw}(k,\delta)$ not exceeding a threshold $\alpha$, i.e. the cumulative distribution function. The $\beta$-Conditional value at risk ($\beta$-CVAR) is:
$\textrm{CVAR}_{\beta}^{zw}(k)=\frac{1}{1-\beta}\int_{L_{zw}(k,\delta)\geq\textrm{VAR}_{\beta}(k)}L_{zw}(k,\delta)p(\delta)d\delta\;.$ (5)
In other words, when looking at the distribution of probability of the loss function $L_{zw}(k,\delta)$, $\text{VAR}_{\beta}^{zw}(k)$ separates the $\beta$ most favorable cases from the $(1-\beta)$ worst cases, while $\textrm{CVAR}_{\beta}^{zw}(k)$ represents the conditional expected value of the loss associated with the worst cases $\delta$ relative to that loss being greater than $\text{VAR}_{\beta}^{zw}(k)$. This is schematically represented in Fig. 2.
Figure 2: Probability density $p^{zw}_{k}(\alpha)$ of the loss function $T_{zw}(\mathrm{s},k,\delta)||$ with $\beta$-VAR, $\beta$-CVAR and deterministic worst case (only defined under Assumption 1).
In this paper, we focus on the problem:
$\begin{array}[]{lll}&\underset{k\in\mathcal{K}}{\text{minimize }}&\textrm{CVAR}_{\beta_{1}}^{z_{1}w_{1}}(k)\\ &\text{subject to }&\textrm{CVAR}^{z_{i}w_{i}}_{\beta_{i}}(k)\leq 1\;,\;2\leq i\leq n\end{array}$ (6)
which replaces problem (2), with the main differences: (i) the parameters $\delta$ are defined as random variables rather than deterministic parameters, over the same bounded support (Assumption 1), (ii) instead of minimizing the worst case without considering any probability, we seek to minimize the conditional expected value of the ($1-\beta$) worst cases.
III Computation of VAR and CVAR
III-A Formulation as an optimization problem
Definition 1 does not allow for a straightforward computation of the VAR and CVAR. Instead, one typically relies on the following result of Rockafellar, in which both are characterized as solutions of a convex optimization problem.
Proposition 1 (Expression of $\beta$-VAR and $\beta$-CVAR as solutions of a convex minimization problem [20])
Under Assumptions 1, 2 and 3, let us define the function
$F^{zw}_{\beta}(k,\alpha)=\alpha+\frac{1}{1-\beta}\int_{\delta\in\mathbb{R}^{m}}[L_{zw}(k,\delta)-\alpha]_{+}p(\delta)d\delta$ (7)
where $x]_{+}=\text{max}(x,0)$. The $\beta$-CVAR of the loss function $L_{zw}(k,\delta)$ is the minimum of $F_{\beta}(k,\alpha)$ as a function of $\alpha$:
$\textrm{CVAR}_{\beta}^{zw}(k)=\min_{\alpha\in\mathbb{R}}F^{zw}_{\beta}(k,\alpha)$ (8)
and the $\beta$-VAR is the left-end value of the set of all $\alpha$ that attain this minimum (possibly reducing to a single point):
$\text{VAR}_{\beta}^{zw}(k)=\text{left end point of }\arg\min_{\alpha\in\mathbb{R}}F^{zw}_{\beta}(k,\alpha)\;.$ (9)
Moreover, $F^{zw}_{\beta}(k,\alpha)$ is convex and continuously differentiable as a function of $\alpha$.
Corollary 1 (Optimal controller for $\beta$-CVAR minimization)
Problem (6) is equivalent to:
$\begin{array}[]{lll}&\underset{k\in\mathcal{K},\alpha\in\mathbb{R}^{n}}{\text{minimize }}&F^{z_{1}w_{1}}_{\beta_{1}}(k,\alpha_{1})\\ &\text{subject to }&F^{z_{i}w_{i}}_{\beta_{i}}(k,\alpha_{i})\leq 1\;,\;2\leq i\leq n\end{array}$ (10)
where we have gathered all $\alpha_{i}$ in a vector $\alpha\in\mathbb{R}^{n}$.
III-B Sample average approximation (SAA)
The function $F^{zw}_{\beta}(k,\alpha)$ can be approximated by sampling NN i.i.d. realizations $\tilde{\delta}^{1}$, $\tilde{\delta}^{2}$, … $\tilde{\delta}^{N}$ from the density $p$ and defining the sample average approximation (SAA) [14]
$\widetilde{F}^{zw}_{\beta}(k,\alpha,\tilde{\delta})=\alpha+\frac{1}{(1-\beta)N}\sum_{i=1}^{N}[L_{zw}(k,\tilde{\delta}^{i})-\alpha]_{+}$ (11)
that converges point-wise to $F^{zw}_{\beta}(k,\alpha)$ when $N\to\infty$ [22].
Optimization of $\beta$-CVAR based on the SAA has been addressed in finance-related literature [20, 7, 26, 25]. As an approximation of the stochastic programming problem (10), we may solve the deterministic SAA problem:
$\begin{array}[]{lll}&\underset{k\in\mathcal{K},\alpha\in\mathbb{R}^{n}}{\text{minimize }}&\widetilde{F}^{z_{1}w_{1}}_{\beta_{1}}(k,\alpha_{1},\tilde{\delta})\\ &\text{subject to }&\widetilde{F}^{z_{i}w_{i}}_{\beta_{i}}(k,\alpha_{i},\tilde{\delta})\leq 1\;,\;2\leq i\leq n\\ &&\underset{\delta\in\mathcal{D}_{\delta}}{\max}\;\lambda^{\mathcal{R}}_{\max}(A(k,\delta))<0\end{array}$ (12)
where $A(k,\delta)$ is the dynamics matrix of the state-space representation of the closed loop, and $\lambda^{\mathcal{R}}_{\max}(\cdot)$ denotes the spectral abscissa. The constraint on the spectral abscissa ensures that Assumption 2 remains satisfied over the whole support $\mathcal{D}_{\delta}$ and not only over the samples $\delta^{(i)}$. How this constraint can be enforced is discussed in Section IV-A.
III-C Interval of confidence
To obtain an accurate solution to the original problem, it is necessary to choose NN sufficiently large. [10, Theorem 3] characterizes the convergence rate and derives the number of samples required to guarantee a prescribed accuracy level with a given confidence. However, this result also depends on the regularity of the uncertain loss function, which is not known a priori. In practice, thanks to the expression of VAR and CVAR as solutions of an optimization problem (Proposition 1), the procedure proposed in [17, Section 3.2] allows to assess the gap between the estimated $\beta$-CVAR:
[…]
and the true one, solution of Eq. (8). Specifically, given a candidate solution $\hat{\alpha}$, one draws $M$ i.i.d. batches of $n$ vectors $\tilde{\delta}^{i1}$, $\tilde{\delta}^{i2}$, … $\tilde{\delta}^{in}$, i=1,2,,Mi=1,2,...,M, and define for each batch:
$G_{i}(n)=\frac{1}{n}\sum_{j=1}^{n}\widetilde{F}^{zw}_{\beta}(k,\hat{\alpha},\tilde{\delta}^{ij})-\min_{\alpha\in\mathbb{R}}\frac{1}{n}\sum_{j=1}^{n}\widetilde{F}^{zw}_{\beta}(k,\alpha,\tilde{\delta}^{ij})\;.$
Since the expected value of the left and right terms are respectively an upper and lower bound of the true optimum, each $G_{i}(n)$ is an observation of a random variable $G(n)$ representing the optimality gap. Noting $\bar{G}(M,n)=\frac{1}{M}\sum_{i=1}^{M}G_{i}(n)$, [17] shows that $\sqrt{M}(\bar{G}(M,n)-E(G(n)))$ converges in distribution to a normal law. Noting $\sigma^{2}_{G}(M,n)$ the sample variance estimator of the variance of $G(n)$, and $t_{M-1,\gamma}$ the $(1-\gamma)$-quantile of a Student $t$ distribution with M1M-1 degrees of freedom, then
$\left[0,\;\bar{G}(M,n)+\frac{t_{M-1,\gamma}\sigma_{G}(M,n)}{\sqrt{M}}\right]$ (13)
is a $(1-\gamma)$-interval of confidence for the optimality gap [17].
IV Optimization
IV-A Optimization procedure
Based on the formulation of previous sections, Algorithm 1 is proposed to tackle the stochastic optimization problem (6).
Algorithm 1 CVAR optimization
Initialization: Active configurations $\mathcal{D}_{\delta}^{a}=\{0\}$, number of samples NN, tolerance $\epsilon$.
Step 1: Find $\hat{k}_{0}$ that robustly stabilizes over the whole support $\mathcal{D}_{\delta}$.
Step 2: Form the SAAs with NN samples, initialize $k=\hat{k}_{0}$ and solve the SAA problem (12) where robust stability over the whole support:
$\underset{\delta\in\mathcal{D}_{\delta}}{\max}\;\lambda^{\mathcal{R}}_{\max}(A(k,\delta))<0$
is replaced with stability over the active configurations:
$\underset{\delta\in\mathcal{D}_{\delta}^{a}}{\max}\;\lambda^{\mathcal{R}}_{\max}(A(k,\delta))<0\;.$
The optimal controller is noted $\hat{k}$.
Step 3: Perform worst-case search in stability:
$\delta^{*}=\textrm{arg }\underset{\delta\in\mathcal{D}_{\delta}}{\max}\;\lambda^{\mathcal{R}}_{\max}(A(\hat{k},\delta))\;.$
If $\lambda^{\mathcal{R}}_{\max}(A(k,\delta^{*}))>0$, add $\delta^{*}$ to $\mathcal{D}_{\delta}^{a}$ and go back to step 2. Otherwise, go to step 4.
Step 4: Evaluate the CVARs of problem (12), with sufficiently tight intervals of confidence. If they are different from those obtained in step 2 by more than a factor $\epsilon$, increase NN and go back to step 2. Otherwise, terminate.
The stability over the whole support $\mathcal{D}_{\delta}$ is addressed through the active configurations approach of [2]; it consists of enforcing stability of configurations belonging to a subset $\mathcal{D}_{\delta}^{a}\subset\mathcal{D}_{\delta}$, initialized with the nominal configuration, and updating this subset through (deterministic) worst-case searches. This procedure is necessary even when the initial controller stabilizes the system over the whole support, because this property may not be maintained when solving the SAA problem (12).
Step 1 allows to provide an initial controller with robust stability over the whole support, as required by Assumption 2. It can be performed with [2], which relies on the algorithm of [4] and the active configurations approach mentioned above; the pure stabilization problem is addressed with the shifted $\mathcal{H}_{\infty}$ norm [9]. One possible, albeit not mandatory, approach is to solve the deterministic problem (2) to obtain an initial controller which also guarantees robust performance with respect to the worst case. The optimization of the SAA approximation is performed in in step 2; it can be addressed with deterministic nonsmooth optimization. The worst-case search for unstable configurations is performed in step 3, as previously discussed. Finally, step 4 evaluates the estimated CVARs obtained with the optimized controller and their intervals of confidence [17] to validate or invalidate the controller.
IV-B Open challenges
While we have proposed a method to address the CVAR optimization, there are still open challenges to address the formulated problem in a more systematic and efficient way.
Future research should focus on developing suitable nonsmooth algorithms to efficiently address step 2 of Algorithm 1, with convergence certificates and able to start from any initial controller (i.e. not requiring step 1). Of particular interest is the work of [18], which proposes an algorithm that tolerates inexact computation of the gradient or of the function itself. This algorithm is specifically designed for lower-$C^{1}$ functions, a class to which the SAA $\widetilde{F}^{zw}_{\beta}(k,\alpha)$ can easily be shown to belong under Assumptions 1, 2 and 3, since this property holds for $T_{zw}(s,k,\delta)||_{\infty}$ as a function of $k$ as long as $T_{zw}(s,k,\delta)$ is internally stable [2, Proposition 2]. In particular, [18] illustrates their method on the $\mathcal{H}_{\infty}$ norm where the exactness of the computation, based on a dichotomy involving the bounded real lemma, can be relaxed. This property is particularly desirable in our SAA where we have to evaluate many times the $\mathcal{H}_{\infty}$ norm, and where our SAA is itself an approximation.
Importance sampling [14, 1] is a technique used in rare event simulation that can reduce the variance of the SAA and reduce the required number of samples. It may particularly be useful when choosing a $\beta$ very close to 1. For example, it was applied to CVAR in [7, 24, 26]. Performing importance sampling in a context of optimization is challenging, since an appropriate sampling distribution may depend on $k$ and therefore has to be updated during optimization [7, 1].
V Application
V-A Benchmark
We consider the benchmark of [15], namely the attitude control of a flexible spacecraft $G(s,\delta)$ to follow a reference $r$. The closed loop is represented in Fig. 3, where $d(s)$ represents actuator dynamics. The control input $u$ of the plant is the torque applied to the satellite and the output vector contains the 3 angle measurements polluted by some noise $n$. We note $k\in\mathcal{K}$ the decision variables describing the controllers $K(s)$ of the desired structure: one 4-th order controller per spacecraft axis, each one taking the measured angle and rate as inputs and generating a control torque. The closed-loop system has 50 states, 43 uncertain parameters $\delta_{i}$, and 72 decision parameters $k_{i}$.
Figure 3: Benchmark: closed-loop system
The uncertain parameters represent mechanical properties. Once normalized, they all follow a Gaussian distribution of mean 0 and standard deviation $1/3$, truncated at 3 $\sigma$, except for one parameter (representing an angle) which follows a uniform distribution between -1 and 1. Additionally, the vectors $\delta$ that do not satisfy a nonlinear constraint $c(\delta)\leq 0$, related to some mechanical constraint detailed in [15], must be removed from the robustness analysis. In the deterministic setting, this constraint can be addressed in the framework of constrained optimization as in [15]. In the $\beta$-CVAR formulation, when performing Monte-Carlo sampling, samples that do not satisfy the constraint are simply discarded; all following results ($\beta$-VAR, $\beta$-CVAR, mean, distribution) refer to the resulting conditional distribution.
The control problem consists in robustly placing a control bandwidth of 0.1 rad/s (Req. 3) while ensuring a modulus margin of 0.5 (Req. 2) and minimizing the variance of the actuator command in response to measurement noise (Req. 1). It involves the 3 loss functions:
[…]
with $W_{1}$ a normalization constant, $W_{2}=\frac{1}{2}$ to ensure the desired modulus margin, $W_{3}(s)=\frac{10\mathrm{s}+1}{2(10\mathrm{s}+0.01)}I_{3}$ a weighting filter to impose the desired bandwidth. Next sections detail the corresponding deterministic and stochastic formulations.
V-B Formulation as a deterministic control problem
In the deterministic formulation, the space of (deterministic) uncertain parameters is the support of $\delta$:
$\mathcal{D}_{\delta}=\{\delta\in[-1,1]^{m}\text{ s.t. }c(\delta)\leq 0\}$
without any consideration of probability. Under the traditional robust control framework, the control problem reads:
$\begin{array}[]{lllr}&\underset{K\in\mathcal{K}}{\text{minimize}}&\underset{\delta\in\mathcal{D}_{\delta}}{\max}\;L_{un}(s,k,\delta)&\text{(Req.~1)}\\ &\text{subject to }&\underset{\delta\in\mathcal{D}_{\delta}}{\max}\;L_{us}(s,k,\delta)\leq 1&\text{(Req.~2)}\\ &&\underset{\delta\in\mathcal{D}_{\delta}}{\max}\;L_{er}(s,k,\delta)\leq 1&\text{(Req.~3)}\end{array}$
and is solved with the method of [15]. Let us note $\hat{k}_{det}$ the decision parameters describing the optimal controller. The resulting densities for $L_{un}(s,\hat{k}_{det},\delta)$ (Req. 1) and $L_{us}(s,\hat{k}_{det},\delta)$ (Req. 2) are presented in Fig. 4 using 10000 samples. It can be verified that the worst case for Req. 2, computed with the deterministic method of [15], is indeed inferior to 1, as imposed. It is also verified for $L_{er}(s,\hat{k}_{det},\delta)$ (Req. 3), not represented for brevity (but the performance metrics are all summarized in Table I). Nominal performance, i.e. when $\delta=0$, is also indicated for completeness.
(a) Density of $L_{un}(s,\hat{k}_{det},\delta)$ (Req. 1)
(b) Density of $L_{us}(s,\hat{k}_{det},\delta)$ (Req. 2)
Figure 4: Empirical probability densities with solution $\hat{k}_{det}$
V-C Formulation as a stochastic control problem
Based on the notations of Definition 1, we fix $\beta=95$% and seek to solve the stochastic problem:
[…]
which we replace, for a set $\tilde{\delta}$ of i.i.d. realizations, by the deterministic problem:
$\begin{array}[]{lllr}&\underset{K\in\mathcal{K},\;\alpha\in\mathbb{R}^{3}}{\text{minimize }}&\widetilde{F}^{un}_{\beta}(k,\alpha_{1},\tilde{\delta})&\text{(Req.~1)}\\ &\text{subject to }&\widetilde{F}^{us}_{\beta}(k,\alpha_{2},\tilde{\delta})\leq 1&\text{(Req.~2)}\\ &&\widetilde{F}^{er}_{\beta}(k,\alpha_{3},\tilde{\delta})\leq 1&\text{(Req.~3)}\\ &&\underset{\delta\in\mathcal{D}_{\delta}}{\max}\;\lambda^{\mathcal{R}}_{\max}(A(k,\delta))<0\end{array}$
We solved this problem with Algorithm 1, using $\hat{k}_{\textrm{det}}$ as initial controller (step 1). Step 2 was performed with Matlab’s function fmincon using the sequential quadratic programming (SQP) algorithm, even though convergence to an optimum is not formally guaranteed due to the nonsmoothness of the SAA. 2500 samples were used initially, and this number was increased to 10000 following step 4. With these 10000 samples, the 99% interval of confidence of Eq. (13) was up to $0.0008$ among the 3 requirements, which was considered sufficient for validation. Stability over the whole domain was also verified with [15]. We note $\hat{k}_{sto}$ the decision parameters describing the optimal controller.
To compare with the results of the deterministic formulation of Section V-B, we also computed the worst case with the method of [15], although this computation was not performed during the optimization itself. The results are presented in Fig. 5 and summarized in Table I. In particular:
1) Fig. 5(b): The $\beta$-CVAR of $L_{us}(s,\hat{k}_{sto},\delta)$ (Req. 2) is 0.98, which is inferior to 1 as imposed. A worst-case of 2.05 (again found with the deterministic method of [15]) is tolerated, whereas it would have been rejected with the traditional formulation. Nevertheless, this worst case is very rare. To illustrate this statement, in addition to the risk measures already presented (VAR, CVAR), we mention that the worst-case performance found among the 1000010000 samples used to draw the histogram is 1.58, and already covers at least 99.9% of the distribution with a confidence of 99.99%. [^2] For a random function $f(\delta)$, noting $\delta^{*}$ the sample realizing the largest function evaluation, [27] shows that $\mathrm{P}(\mathrm{P}(f(\delta)>f(\delta^{*})\leq\epsilon)\geq 1-\gamma$ is guaranteed when NN verifies: $N>\frac{\textrm{ln}(\gamma)}{\textrm{ln}(1-\epsilon)}\;.$
2) Fig. 5(a): As a result of relaxing Req. 2 (from worst case to CVAR criterion), the Req. 1 is considerably improved: all metrics were reduced by $\sim$30%.
(a) Density of $L_{un}(s,\hat{k}_{sto},\delta)$ (Req. 1)
(b) Density of $L_{us}(s,\hat{k}_{sto},\delta)$ (Req. 2)
Figure 5: Empirical probability densities with solution $\hat{k}_{sto}$
TABLE I: Performance metrics for the deterministic (det.) and stochastic (sto.) approaches.
Req.1 Nominal Mean $\beta$-VAR $\beta$-CVAR Worst
det. 1.02 1.03 1.09 1.10 1.17
sto. 0.69 0.70 0.73 0.74 0.79
Req.2
det. 0.90 0.87 0.91 0.92 0.93
sto. 0.87 0.79 0.87 0.98 2.05
Req.3
det. 1.00 1.00 1.00 1.00 1.00
sto. 1.00 1.00 1.00 1.00 1.05
VI Conclusion
To reduce the conservatism of conventional robust $\mathcal{H}_{\infty}$ control, which minimizes a worst-case performance that may be very unlikely, we introduced an alternative stochastic approach based on the conditional value at risk. Its capacity to significantly improve the performance of uncertain control systems, at the expense of tolerating rare degrading worst-case configurations, was illustrated on a mechanical system.
ACKNOWLEDGMENT
The authors acknowledge the use of an AI language model (ChatGPT: GPT-5.2) to assist in improving the wording and reformulation of selected sentences.